OpenStudy (anonymous):

Prove: cos3x/cosx=1-4sin^2x

OpenStudy (anonymous):

I got up to cos3xsecx=1-4sin^2x

zepdrix (zepdrix):

Mmm no we don't want that secant ^^ We want to apply the Sum Angle Formula for Cosine to the top, And then the Double Angle Formula for Cosine after that. \[\cos3x=\cos(2x+x)\] Recall our Sum Angle Formula for Cosine:\[\large \cos(\alpha+\beta)=\cos \alpha \cos \beta-\sin \alpha \sin \beta\]

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

What would I do after that @zepdrix ?

OpenStudy (anonymous):

\[\cos2x \cos x-\sin2xsinx=cosx-4\sin^2xcosx\]

zepdrix (zepdrix):

Ok looks good so far! From here, let's factor a cosx out of each term.

OpenStudy (anonymous):

How would I factor a cosx out of a sinx?

OpenStudy (anonymous):

Or did you just mean the cos?

zepdrix (zepdrix):

The cosine is not `inside` of the sine function, it's being multiplied by it. Let \(\large a=\cos x\) Then we have \(\large a-4a \sin^2x\) Factor an a of out of each term :)

OpenStudy (anonymous):

cos3x = cos(x) cos (2x) - sin(x) sin (2x) = cos(x) (cos² (x) - sin²(x)) - sin(x) (sinx cosx +cos x sin x) = cos(x) (1 - sin² (x) - sin²(x)) - 2 sin² x cos x = cos x (1 - 4 sin² x) Dividing by cos x gives: cos 3x / cos x = 1 - 4 sin² x Formulae we used in the calculation: cos (a+b) = cos a cos b - sin a sin b sin (a+b) = sin a cos b + cos a sin b cos² a = 1 - sin² a

OpenStudy (anonymous):

I mean out of the left side. The right side I understand.

zepdrix (zepdrix):

Yes I meant the left side also :o

OpenStudy (anonymous):

Where is the sin being multiplied by cos on the left side?

zepdrix (zepdrix):

\[\large \cos x-4\sin^2x(\cos x) \qquad = \qquad \cos x-4(\cos x)\sin^2x\] Multiplication is commutative, so you can do it in any order. So I just moved the cosine over, so it's a little easier to read.

OpenStudy (anonymous):

I'm talking about the other side of the equation (cos2xcosx-sin2xsinx)

zepdrix (zepdrix):

Oohhhh I see what happened, you multiplied the cosine over to the right? Ok ok ok I misunderstood, I thought you worked out the double angle formulas already. If you're trying to show an identity is true, do not do anything to the other side, we want to keep that cosx in the denominator. Ok ok, let's see.

zepdrix (zepdrix):

We want to apply these two formulas:\[\large \cos2x=1-2\sin^2x\] \[\large \sin2x=2\sin x \cos x\]

OpenStudy (anonymous):

Wait, so we don't do anything to the cos3x/cosx?

zepdrix (zepdrix):

We can't move any of it to the right side.

OpenStudy (anonymous):

I thought we could multiply by cosx on both sides.

zepdrix (zepdrix):

No :o

OpenStudy (anonymous):

So we have cos3x=cosx-4sinx^2cosx

zepdrix (zepdrix):

If you're "proving" something, it means you start with one side, and end with the other side. You don't manipulate both sides and meet them half way.

OpenStudy (anonymous):

Wouldn't the results be the same either way?

OpenStudy (anonymous):

Tehcnically speaking.

zepdrix (zepdrix):

It would, yes :) But the correct process of working through proofs is kind of important.

OpenStudy (anonymous):

Oh, then everything I know is a lie. :P

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

Okay then, so we leave it at cos3x/cosx?

zepdrix (zepdrix):

Yes :) Don't move anything to the right. We have to use trig identities to `change it`. So here's what we've done so far. \[\large \frac{\cos3x}{\cos x}=1-4\sin^2x\] \[\large \frac{\cos2x \cos x-\sin2x \sin x}{\cos x}=1-4\sin^2x\]

zepdrix (zepdrix):

\[\large \color{royalblue}{\cos2x=1-2\sin^2x}\]\[\large \color{purple}{\sin2x=2\sin x \cos x}\]We're going to use these two identities to change things around.\[\large \frac{\color{royalblue}{\cos2x} \cos x-\color{purple}{\sin2x} \sin x}{\cos x}=1-4\sin^2x\] Understand what to do next? :o

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

Okay, so it becomes 1-2sin^2xcosx-2sinxcosxsinx=1-4in^2x which then becomes 1-2sin^2xcosx-3sinxcosx=1-4sin^2x

OpenStudy (anonymous):

Right?

zepdrix (zepdrix):

Mmm your math looks like a little funky. Should use brackets when we apply this first identity,\[\large \frac{(\color{royalblue}{1-2\sin^2x}) \cos x-\color{purple}{2\sin x \cos x} \sin x}{\cos x}=1-4\sin^2x\]

OpenStudy (anonymous):

Okay, it looks better on paper, :P

OpenStudy (anonymous):

So we are left with 1-2sin^2x-2sin^2x?

zepdrix (zepdrix):

Oh you divided out the cosine from the denominator? Yah sounds good!

OpenStudy (anonymous):

Okay, just making sure, and is my end correct?

OpenStudy (anonymous):

Or would it be 3sin instead of 2sin^2x?

OpenStudy (anonymous):

The second one, not the frist 2sin^2x.

zepdrix (zepdrix):

So it looks like you got it to this point: \[\large 1-2\sin^2x-2\sin^2x \qquad=\qquad 1-4\sin^2x\] Which is correct. Understand what to do next? :)

zepdrix (zepdrix):

\[\large 2\sin x \cos x \sin x \qquad=\qquad 2\sin^2x \cos x\]Just in case there was any confusion on that part.

OpenStudy (anonymous):

Wait, I don't get it.

zepdrix (zepdrix):

Don't get what? That last multiplication step?

OpenStudy (anonymous):

What did you multiply? Sorry, I had to do the dishes.